Ahmer walks 795m west and 55m north 80m east find the displacement and distance?
1
Expert's answer
2020-03-09T10:55:49-0400

The distance


"d=795\:\rm m+55\:\rm m+80\:\rm m=930\:\rm m."

The displacement


"s=\sqrt{(795-80)^2+55^2}=717\:\rm m."